### v2

@@ Abstract.
[…]
@@ 1. Introduction
In optimization theory, when studying the limit of a sequence of optimization problems, one seeks a natural notion of convergence of the data that guarantees convergence of both optimal values and minimizers; in other words, a form of variational convergence that allows one to pass to the limit in variational problems. The first concept of this kind was introduced by Wijsman in finite dimensions under the name infimal convergence [78, 79], motivated by questions in statistical decision theory [31]. In the 1970s, De Giorgi and Franzoni extended this perspective to general topological spaces under the name $\Gamma$-convergence [47], which has since become one of the central tools of the calculus of variations.
In parallel, the infinite-dimensional setting introduced new challenges due to the role of weak topologies. To address these, Mosco introduced in the late 1960s a convergence notion for lower semicontinuous convex functions—and hence for closed convex sets via their indicator functions—in order to study the stability of solutions to variational inequalities [55, 57, 56]. This notion, now known as Mosco convergence, was closely linked to the convergence of solutions of elliptic and parabolic PDEs and the associated energy functionals (see, e.g., [75, 40, 74]).
A natural line of investigation is how $\Gamma$-convergence and Mosco convergence interact with fundamental operations in convex analysis and optimization, such as Fenchel conjugation, polarity of convex sets, or subdifferentiation. The continuity of the Fenchel conjugate with respect to variational convergences was first analyzed by Wijsman [79], who proved its continuity under infimal convergence. Around the same time, Walkup and Wets [77] showed that in reflexive Banach spaces the polarity operation for closed convex cones is an isometry; as Rockafellar later noted [66, Theorem 14.4], this already implies continuity of the Fenchel transform in finite dimensions. Related connections between polarity and convex cones were further developed in [65].
The breakthrough in infinite dimensions came with Mosco’s 1971 theorem [58], establishing that Mosco convergence makes the Fenchel conjugate bicontinuous. Reflexivity itself can be characterized by this bicontinuity property [20, 18]. Mosco’s theorem thus extends Wijsman’s results and ensures continuity of conjugation also in finite dimensions. Subsequent work by Joly [48] generalized Mosco’s results within Banach spaces, while McLinden and Bergstrom [54] explored consequences in finite dimensions for convex sets, functions, and programming models. In the following decades, continuity of the Fenchel conjugate and the polar map, and generalizations of Mosco convergence to non-reflexive spaces, as well as topologies for closed (or closed and convex) sets and convex functions, were extensively studied (see, e.g., [15, 16, 17, 7, 42, 8, 35, 60, 24, 73, 23, 22, 25, 27, 21, 19, 61]), with Beer and Borwein’s result [20] showing that bicontinuity of the Fenchel conjugate is in fact equivalent to reflexivity.
@@
A further milestone was reached by Attouch in 1977, who proved that Mosco convergence of convex functions is equivalent to the convergence of their subdifferentials in the sense of Painlevé–Kuratowski set convergence, up to a normalization constant [3]. This fundamental result, now known as Attouch theorem, has become a cornerstone of variational analysis. It’s applications are numerous: to stability in optimization and variational problems [14, 13, 9, 10, 11, 12], to the analysis of parabolic variational inequalities [6], to numerical stability in optimization algorithms [51], and to structural results such as generalized second-order derivatives of convex functions [68, 67, 34, 33, 52, 63] and differentiability of Lipschitz set-valued mappings [39]. Attouch’s theorem also remains influential in applied areas such as nonsmooth mechanics and the study of subgradient evolution systems [5].
Today, $\Gamma$-convergence is well-recognized as a unifying framework across diverse fields of pure and applied mathematics. While its origins lie in mathematical programming, it has become an indispensable tool in the approximation of optimization problems, the study of subdifferential inclusions, consistency in stochastic programming, calculus of variations, PDEs arising as Euler–Lagrange equations of variational problems, and even extensions of probabilistic principles such as the strong law of large numbers for functions (see, e.g., [2, 69, 70, 49, 72, 38, 71, 29, 5, 4] and the references therein).
@@ 3.1. Notations and basic definitions
Finally, $X$ is said to be a weakly compactly generated (WCG) space if there exists a weakly compact set $A\subset X$ such that the linear span of $A$ is dense in $X$ with respect to the norm topology. In this definition, $A$ may be assumed, without loss of generality, to be convex and symmetric (see, e.g., [43, Chapter 13]). For some examples of $WCG$ spaces, we refer the reader to [43, Chapter 13 and Theorem 14.9].
@@
Furthermore, a function $f\colon X\to\mathbb{R}\cup\{\infty\}$ will be called proper if ${\rm dom}\kern 1.19995ptf\neq\emptyset$, and will be called convex if ${\rm epi}\kern 1.19995ptf$ is a convex set.
@@ 3.2. Limits of sequences of functions and operators
where $x_{n}\overset{\tau}{\to}x$ denotes convergence of the sequence $(x_{n})_{n\in\mathbb{N}}$ with respect to the topology $\tau$. Moreover, it is a known fact that when the topology $\tau$ is metrizable (or more general, first countable), then each infimum is indeed a minimum [29, p.30, Remark 1.26].
@@ 4. Limits of convex functions and subdifferentials
In this section we study the relation between the epigraphical convergence of a sequence of convex functions and the $w^{\ast}$-sequential limit of its conjugate functions, in the sense of (3.1).
The first result of this section provides a characterization of the conjugate of the lower sequential epigraphical limit of a sequence of functions, as the upper sequential epigraphical limit of their corresponding conjugates with respect to the weak-topology. The proof of this result will be presented in Section 6.
@@
These sets can be regarded as subsets of $X^{\ast}$, using the identification $\ell_{1}(\mathbb{N})\subset X^{\ast}=\ell_{\infty}(\mathbb{N})^{\ast}=\left(\ell_{1}(\mathbb{N})\right)^{\ast\ast}$. It is not difficult to show that each $C_{k}$ is $w^{\ast}$-closed and convex. Furthermore, by Goldstine’s theorem (see, e.g., [43, Theorem 3.96]), we obtain
@@
Finally, by the Odell–Rosenthal theorem (see, e.g., [59] or [41, Chapter XIII, Theorem 10]), the closed unit ball in $\ell_{1}(\mathbb{N})$ is not weak sequentially dense in the closed unit ball of $X^{\ast}$. Particularly, it implies that
@@
By taking the Fenchel conjugate in (4.2) (see, e.g., [28, Fact 4.4.9]), one obtains that this condition is equivalent to the uniform boundedness of the functions in a neighborhood of the origin, which means, there exist $\rho,\varepsilon>0$ such that
@@
Furthermore, by [81, Theorem 2.2.22], this is in turn equivalent to the existence of a convex open neighborhood $U$ of the origin such that, for every $x\in U$, the sequence $(f_{n}(x))_{n\in\mathbb{N}}$ is bounded.
Now, let us recall the following result established first by Mosco in [57] and generalized by Joly in [48]. Nonetheless, a proof can be found in [62, Theorem 2].
@@
As an immediate consequence of the two preceding results, we obtain a dual formula that relates the epigraphical limit in the strong topology to the epigraphical limits of the corresponding conjugates with respect to the weak topology. The proof of this result is presented in Section 7.
@@
Next, we establish a connection between the epigraphical limits of a sequence of convex functions and the convergence of their associated subdifferentials in the sense of (3.2). It is worth noting that the $\|\cdot\|\times w^{\ast}$ sequential convergence of subdifferentials has already been investigated in [80]. More precisely, in [80, Theorem 3.1] the authors proved that if a sequence of proper, lower semicontinuous convex functions Mosco-converges to $f$ and satisfies (4.3), then the corresponding subdifferentials converge with respect to $\|\cdot\|\times w^{\ast}$ in the sense of (3.2).
In what follows, we show that this implication can be strengthened: the uniform boundedness condition in (4.3) is in fact equivalent to the epigraphical convergence of the functions. The proof of this equivalence is deferred to Section 8.
@@
Observe that [80, Example 3.2] also serves as a counterexample to Theorem 4.7, showing that $(c)$ does not imply $(b)$ once the WCG hypothesis is dropped.
In the setting of a reflexive Banach space $X$, the roles of $X$ and $X^{\ast}$ can be interchanged in the above result. Hence, the equicoercivity assumption (4.1) can be formulated in the primal space. Moreover, since in this case the $w^{\ast}$-topology on $X^{\ast}$ coincides with the weak topology, we obtain the following corollary as a direct consequence of Theorem 4.7.
@@
Lastly, we get the following result, which relates the $G(\left\lVert\cdot\right\rVert\times w^{\ast})$, the $G(w\times\left\lVert\cdot\right\rVert_{\ast})$ and the $G(\left\lVert\cdot\right\rVert\times\left\lVert\cdot\right\rVert_{\ast})$ convergence of subdifferentials in the reflexive case.
@@
If the sequence $(f_{n})_{n\in\mathbb{N}}$ satisfies (4.4), then
$(a)\wedge(b)\Longleftrightarrow(c).$

### v3

@@ Abstract.
We generalize the duality principle for the $\Gamma$-convergence of proper convex lower semicontinuous functions, which was previously known only in separable reflexive Banach spaces. Aditionally, We extend this duality to the broader class of weakly compactly generated (WCG) Banach spaces. Under a suitable hypotheses, we show that $\Gamma$-convergence in the strong topology is equivalent to $\Gamma$-convergence of the Fenchel conjugates in the weak topology. We further prove that this duality is equivalent to the graphical convergence of the associated subdifferentials with respect to the product topology given by the norm on the primal space and the weak topology on the dual. The extension to the WCG framework allows the inclusion of every separable space and every reflexive space, as well as important non-reflexive that might not be separable, such as $L^{1}(\mu)$ for an arbitrary $\sigma$-finite measure. As an application, we derive dual characterizations of the $\Gamma$-convergence of convex integral functionals on $L^{p}$ spaces. The proof of the main result relies on a separable reduction approach in WCG spaces using modern tools of Banach space theory, which, as a byproduct, provides new insight into the internal structure of WCG Banach spaces, in a spirit closely related to the structural phenomena underlying the theory of Asplund spaces [Cúth and Fabian (2016)].
@@ 1. Introduction
In optimization and the calculus of variations, one seeks a notion of convergence for a sequence of variational problems that ensures the convergence of both minimizers and minimal values. The first concept of this kind was introduced by Wijsman in finite dimensions under the name infimal convergence [76, 77], motivated by questions in statistical decision theory [32]. The next step came from variational inequalities (see [55]), when Mosco introduced in the late 1960s a convergence notion in reflexive spaces for lower semicontinuous convex functions—and hence for closed convex sets via their indicator functions—in order to study the stability of solutions to variational inequalities [59, 61, 60].This notion, now known as Mosco convergence, was closely related to the convergence of solutions of elliptic and parabolic PDEs and the associated energy functionals (see, e.g., [74, 41, 73]).In the 1970s, De Giorgi and Franzoni extended this perspective to general topological spaces and general functions under the name $\Gamma$-convergence [49], which has since become one of the central tools of the calculus of variations (see, for e.g, [39, 30]).
The dualization of variational convergences for convex functions was initiated by Wijsman in finite dimension in [77]. It was continued and generalized by Mosco to the reflexive setting in [62], where he established that Mosco convergence renders the Fenchel conjugation bicontinuous. This phenomenon is, in fact, characteristic of reflexivity (see [16, 23]). Subsequently, Joly extended Mosco’s results to general Banach spaces (see [50]. For $\Gamma$-convergence itself, a duality was established in reflexive separable Banach spaces: Provided a boundness assumption in the primal, a sequence of convex functions $(f_{n})_{n\in\mathbb{N}}$ (sequentially) $\Gamma$-converges with the weak topology to a function $f$ if and only if the sequence of conjugates $\Gamma$-converges with the norm topology to the conjugate of $f$, as well as applications to convex integrals of the calculus of variations (see [57, 25, 56] and references therein). On the other hand, McLinden and Bergstrom [58] explored consequences in finite dimensions for convex sets, functions, and programming models.
@@
A further milestone in the dualization of $\Gamma$-convergence was reached by Attouch in 1977, who proved that Mosco convergence of convex functions is equivalent to the graphical convergence of their subdifferentials up to a normalization constant [10]. This fundamental result, now known as Attouch theorem, has become a cornerstone of variational analysis. It’s applications are numerous: to stability in optimization and variational problems [14, 13, 6, 7, 8, 9], to the analysis of parabolic variational inequalities [3], to numerical stability in optimization algorithms [52], and to structural results such as generalized second-order derivatives of convex functions [69, 68, 34, 35, 53, 67] and differentiability of Lipschitz set-valued mappings [40]. Attouch’s theorem also remains influential in applied areas such as nonsmooth mechanics and the study of subgradient evolution systems [2].
Despite these developments, the duality principle for $\Gamma$-convergence has remained confined, to the best of our knowledge, to reflexive separable Banach spaces. The role of reflexivity and separability in this context is not merely technical: both enter naturally in the classical proofs through compactness and metrizability arguments. It is therefore natural to ask whether these assumptions are intrinsic to the duality mechanism itself.
The main result of this paper shows that they are not, and that the natural dual framework for $\Gamma$-convergence of convex functions is given by the strong topology in the primal space and the weak topology in the dual. Within this framework, we establish a Fenchel duality principle for $\Gamma$-convergence in weakly compactly generated Banach spaces.
Our results also provide an answer to a question raised by Fitzpatrick and Lewis in [48], concerning the relation between the convergence of convex functions and the convergence of their Fenchel conjugates, moreover, the result covered by [48] is a particular case of our main result. Fitzpatrick question raised at a time where continuity of the Fenchel conjugate and the polar map, and generalizations of Mosco convergence to non-reflexive spaces, as well as topologies for closed (or closed and convex) sets and convex functions, were extensively studied (see, e.g., [15, 21, 22, 4, 43, 5, 36, 65, 20, 72, 19, 18, 26, 28, 17, 24, 66]).
@@
The proof of our main result relies heavily on modern tools of Banach Space Theory. Our approach also reveals insights on the structural properties of weakly compactly generated Banach spaces, analogous to those used in the recent characterization of Asplund spaces by Fabian and Cúth in [38].
@@ 3.1. Notations and basic definitions
The Banach space $X$ is said to be a weakly compactly generated (WCG) space if there exists a weakly compact set $A\subset X$ such that the linear span of $A$ is dense in $X$ with respect to the norm topology. In this definition, $A$ may be assumed, without loss of generality, to be convex and symmetric (see, e.g., [44, Chapter 13]). We recall that every separable or reflexive space is WCG. There are also examples of WCG spaces that are not reflexive and not necessarily separable, such as $L^{1}(\mu)$ for any $\sigma$-finite measure $\mu$, as well as the $C(K)$ space whenever $K$ is a compact topological space homeomorphic to a weakly compact subset of a Banach space endowed with the restriction of the weak topology (see [44, Chapter 13 and Theorem 14.9]).
@@
Furthermore, a function $f\colon X\to\mathbb{R}\cup\{\infty\}$ will be called proper if ${\rm dom}\kern 1.19995ptf\neq\emptyset$, it will be called convex if ${\rm epi}\kern 1.19995ptf$ is a convex set, and will be called lower semicontinuous if ${\rm epi}\kern 1.19995ptf$ is closed . We can also define the
@@ 3.2. Limits of sequences of functions and operators
where $x_{n}\overset{\tau}{\to}x$ denotes convergence of the sequence $(x_{n})_{n\in\mathbb{N}}$ with respect to the topology $\tau$. Moreover, it is a known fact that when the topology $\tau$ is metrizable (or more general, first countable), then each infimum is indeed a minimum [30, p.30, Remark 1.26].
@@ 4. Limits of convex functions and subdifferentials
In this section we study the relation between the $\Gamma$-convergence of a sequence of convex functions and the $w^{\ast}$-sequential limit of its conjugate functions, in the sense of (3.1).
[…]
@@
These sets can be regarded as subsets of $X^{\ast}$, using the identification $\ell_{1}(\mathbb{N})\subset X^{\ast}=\ell_{\infty}(\mathbb{N})^{\ast}=\left(\ell_{1}(\mathbb{N})\right)^{\ast\ast}$. It is not difficult to show that each $C_{k}$ is $w^{\ast}$-closed and convex. Furthermore, by Goldstine’s theorem (see, e.g., [44, Theorem 3.96]), we obtain
@@
Finally, by the Odell–Rosenthal theorem (see, e.g., [63] or [42, Chapter XIII, Theorem 10]), the closed unit ball in $\ell_{1}(\mathbb{N})$ is not weak sequentially dense in the closed unit ball of $X^{\ast}$. Particularly, it implies that
@@
By taking the Fenchel conjugate in (4.2) (see, e.g., [29, Fact 4.4.9]), one obtains that this condition is equivalent to the uniform boundedness of the functions in a neighborhood of the origin, which means, there exist $\rho,\varepsilon>0$ such that
@@
Furthermore, by [79, Theorem 2.2.22], this is in turn equivalent to the existence of a convex open neighborhood $U$ of the origin such that, for every $x\in U$, the sequence $(f_{n}(x))_{n\in\mathbb{N}}$ is bounded. Moreover, in this last case, i.e, when condition (4.3) is satisfied, the $\Gamma$-limit of the sequence $(f_{n})_{n\in\mathbb{N}}$ coincides with its pointwise limit in $\mathbb{B}_{\varepsilon}(0)$ (see [39, Proposition 5.12]).
Now, let us recall the following result provided in [12][Theorem 7.5.1] and attributed to Mosco (see also [64, Theorem 2]).
@@
As an immediate consequence of the two preceding results, we obtain a dual formula that relates the $\Gamma$-limit in the strong topology to the $\Gamma$-limit of the corresponding conjugates with respect to the weak topology. The proof of this result is presented in Section 8.
@@
Next, we establish a connection between the $\Gamma$-limit of a sequence of convex functions and the graphical convergence of the subdifferentials (see (3.2)). It is worth noting that the $\|\cdot\|\times w^{\ast}$ sequential convergence of subdifferentials has already been investigated in [78]. More precisely, in [78, Theorem 3.1] the authors proved that if a sequence of proper, lower semicontinuous convex functions Mosco-converges to $f$ and satisfies (4.3), then the corresponding subdifferentials converge with respect to $\|\cdot\|\times w^{\ast}$ in the sense of (3.2).
In what follows, we show that this implication can be strengthened and is sharper. The proof of this equivalence is deferred to Section 9.
@@
Observe that [78, Example 3.2] also serves as a counterexample to Theorem 4.7, showing that $(c)$ does not imply $(b)$ once the WCG hypothesis is dropped.
In the case of a locally equibounded sequence of functions on a normed space, $\Gamma$-convergence coincides with pointwise convergence; hence, we deduce the following corollary.
Corollary 4.8.
Let $X$ be a WCG space, and let $f,f_{n}\colon X\colon\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. In addition, suppose that $(f_{n})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $x\in X$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall y\in\mathbb{B}_{r}(x),\;|f_{n}(y)|\leq M.$
Then, the following statements are equivalent:
  1. (a)
    $f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}$.
  2. (b)
  3. (c)
    $f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}$.
  4. (d)
    For every $x\in X$, $\lim_{n\to\infty}f_{n}(x)=f(x)$.
[…]
In the setting of a reflexive Banach space $X$, the roles of $X$ and $X^{\ast}$ can be interchanged in the above result. Hence, the equicoercivity assumption (4.1) can be formulated in the primal space. Moreover, since in this case the $w^{\ast}$-topology on $X^{\ast}$ coincides with the weak topology, we obtain the following corollaries as a direct consequence of Theorem 4.7 and Corollary 4.8, respectively.
@@
Corollary 4.10.
Let $X$ be a reflexive space, and let $f,f_{n}\colon X\colon\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. In addition, suppose that $(f^{\ast}_{n})_{n\in\mathbb{N}}$ is locally equibounded, i.e, for every $x^{\ast}\in X^{\ast}$, there exists $r>0$ and $M\in\mathbb{R}$ such that
$\forall n\in\mathbb{N},\forall y^{\ast}\in\mathbb{B}^{\ast}_{r}(x^{\ast}),\;|f^{\ast}_{n}(y^{\ast})|\leq M,$
where $\mathbb{B}^{\ast}_{r}(x^{\ast})$ denotes the open ball of center $x^{\ast}$ and radius $r$ with the norm $\left\lVert\cdot\right\rVert_{\ast}$. Then, the following statements are equivalent:
  1. (a)
    $f={\Gamma(w)\text{-}\displaystyle\lim_{n}}f_{n}$.
  2. (b)
  3. (c)
    $f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}$.
  4. (d)
    For every $x^{\ast}\in X^{\ast}$, $\lim_{n\to\infty}f^{\ast}_{n}(x^{\ast})=f^{\ast}(x^{\ast})$.
We also get the following corollary, which relates the $G(\left\lVert\cdot\right\rVert\times w^{\ast})$, the $G(w\times\left\lVert\cdot\right\rVert_{\ast})$ and the $G(\left\lVert\cdot\right\rVert\times\left\lVert\cdot\right\rVert_{\ast})$ convergence of subdifferentials in the reflexive case.
@@
  4. (d)
    $f^{\ast}=\Gamma\text{-}\displaystyle\lim_{n}f^{\ast}_{n}$ and $f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}$.
If the sequence $(f_{n})_{n\in\mathbb{N}}$ satisfies (4.4) or if the sequence $(f^{\ast}_{n})_{n\in\mathbb{N}}$ satisfies (4.1), then
$((a)\wedge(b))\Longleftrightarrow(c)\Longleftrightarrow(d).$
Proof. That $(c)$ implies $(a)$ and $(b)$ follows directly from [11, Proposition 3.59]. That $(c)$ implies $(d)$ follows from the Attouch Theorem and the fact that Mosco convergence makes the Fenchel conjugate bicontinuous in reflexive spaces. That $(a)$ and $(b)$ implies $(d)$ follows from Theorem 4.7 and Corollary 4.9. Now, assume that the sequence $(f_{n})_{n\in\mathbb{N}}$ satisfies (4.4). Suppose that $(a)$ and $(b)$ hold. Using Theorem 4.7 and Corollary 4.9, we deduce that $f_{n}$ Mosco-converges to $f$. Hence, using the Attouch theorem, we conclude $(c)$. To see that $(d)$ implies $(c)$ (asumming that $(f_{n})_{n\in\mathbb{N}}$ satisfies (4.4)), we use Corollary 4.9 to deduce that $f={\Gamma(w)\text{-}\displaystyle\lim_{n}}f_{n}$, and then $f_{n}$ Mosco converges to $f$, so we conclude $(c)$ by the Attouch theorem.
In order to see that $(d)$ implies $(c)$ when $(f^{\ast}_{n})_{n\in\mathbb{N}}$ satisfies 4.1, we use Theorem 4.7 to deduce that $f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}$, and therefore, $f^{\ast}_{n}$ Mosco converges to $f^{\ast}$, hence, using that the Fenchel conjugate is bicontinuous with respect to Mosco Convergence, we obtain that $f_{n}$ Mosco converges to $f$. Thence, we conclude $(c)$ using the Attouch theorem.